# Robust Containment Queries over Collections of Trimmed NURBS Surfaces via Generalized Winding Numbers

### v2

@@ Abstract.
We propose a containment query that is robust to the watertightness of regions bound by trimmed NURBS surfaces, as this property is difficult to guarantee for in-the-wild CAD models. Containment is determined through the generalized winding number (GWN), a mathematical construction that is indifferent to the arrangement of surfaces in the shape. Applying contemporary techniques for the 3D GWN to trimmed NURBS surfaces requires some form of geometric discretization, introducing computational inefficiency to the algorithm and even risking containment misclassifications near the surface. In contrast, our proposed method uses a novel reformulation of the relevant surface integral based on Stokes’ theorem, which operates on the boundary and trimming curves as provided through rapidly converging adaptive quadrature. Batches of queries are further accelerated by memoizing (i.e. caching and reusing) quadrature node positions and tangents as they are evaluated. We demonstrate that our GWN method is robust to complex trimming geometry in a CAD model, and is accurate up to arbitrary precision at arbitrary distances from the surface. The derived containment query is therefore robust to model non-watertightness while respecting all curved features of the input shape.
@@ 1. Introduction
In computer graphics and adjacent fields, 3D objects are often defined in terms of a boundary representation (B-Rep) of the surface. While B-reps are very useful for flexibly designing 3D objects within a CAD system, they often fail to be watertight, in the sense that they do not define a legitimate interior volume (see Figure 2). For example, disconnected edges between component surfaces may be an intentional design decision, but are just as likely to result from human error during modeling, or differing tolerances between applications (Kasik et al., 2005; McAdams et al., 2011) (see Figure 2). The use of Boolean operations during model construction can also introduce problematic non-manifold edges. Issues of watertightness are greatly exacerbated in the context of trimmed NURBS (non-uniform rational B-spline) surfaces, where such gaps are nearly mathematically inevitable due to the complexities of projecting 3D surface-surface intersection curves into 2D parametric trimming curves (Sederberg et al., 2008; Zou et al., 2024). Nevertheless, we are often interested in efficiently and robustly determining whether a given point is contained within the represented object.
@@
Similarly, methods for immersogeometric analysis model fluid-structure interactions are by placing a complicated CAD model on a simpler computational grid which records the numerical solution (Kamensky et al., 2015; Fromm et al., 2023). In such methods, it is necessary to know whether points on this grid are inside or outside the CAD model, which can be particularly challenging for points located arbitrarily close to the B-Rep.
@@
  • We extend the theory of generalized winding numbers to the context of trimmed NURBS patches. With this, we define a containment query for B-Reps composed of unstructured collections of such shapes that is robust to, and indeed altogether indifferent to, watertightness.
  • We describe a novel algorithmic framework for evaluating the GWN with respect to a trimmed NURBS surface through Stokes’ theorem, greatly improving efficiency and stability by reformulating the relevant surface integral as a line integral along the surface boundary. We evaluate these integrals with an adaptive quadrature scheme to achieve results up to arbitrary accuracy, and further improve performance with a memoization strategy that caches and reuses geometric data for each surface. Within our algorithm, we robustly handle trim tests (visibility queries over trimming curves in the 2D parameter space of each surface) with the 2D GWN.
@@ 2.1. Containment Queries in 3D CAD Applications
There are many strategies to evaluate containment queries for CAD applications. In the typical ray-casting method, an arbitrary ray is extended out from the point of interest and the number of (oriented) intersections with bounding geometry are recorded. Assuming a watertight B-Rep, each such intersection increments a crossing number whose parity determines containment.
@@
There is considerable overlap with the adjacent task of ray-tracing for the purposes of rendering, which typically considers a specific, pre-selected ray direction, and only requires calculation of the first intersection of a ray with the B-Rep. Much of the research focus in ray-tracing relates to triangulated surfaces, as many hardware platforms are capable of calculating ray-triangle intersections exactly and massively in parallel (Haines and Akenine-Möller, 2019, Chapter 3). However, there is also interest in applying such techniques directly to trimmed NURBS surfaces (Schollmeyer and Fröhlich, 2009, 2019; Xiong et al., 2023; Zhu et al., 2025). Of particular note are recent methods for ray-tracing bilinear patches, as, similar to triangles, they admit direct geometric formulae for computing their intersections with rays (Reshetov, 2019).
On more general curved surfaces, the surface-line intersection routines used to determine containment are typically more complex and iterative. For example, algebraic methods find solutions in terms of the parametrization of the ray and/or surface, either through numerical root-finding (Martin et al., 2000), convex optimization (Marschner et al., 2021), or even via singular value decompositions of matrix embedding problems (Shen et al., 2016). These algebraic methods can find intersection points to a high degree of numerical precision, but at the cost of additional sensitivity to the initial configuration of the relevant solver. On the other hand, geometric subdivision methods successively clip the surface at points where intersections are not possible, resulting in useful guarantees for their accuracy (Nishita et al., 1990; Efremov et al., 2005).
Regardless of the underlying method, ray-casting methods are highly sensitive to the specific configuration of surface and cast ray, even if the direction of the ray is arbitrary in principle. For example, accurately handling rays which are tangent to the B-Rep or intersect the boundary of a component surface requires additional processing and branching. These issues are compounded in the non-watertight case, where a cast ray passing through a numerical gap between surfaces results in an incorrect and unexpected containment result. There exist methods of mitigating these issues, such as by casting multiple rays and averaging the results (Nooruddin and Turk, 2003), but winding number methods avoid them entirely.
The integral winding number has been long understood as an alternative to ray-casting for closed, linear shapes in both 2D (Hormann and Agathos, 2001) and 3D (Carvalho and Cavalcanti, 1995). Because the result is independent of any particular direction, these methods implicitly handle configurations which are otherwise difficult for ray-casting. At the same time, the winding number is also indifferent to watertightness, as the signed angle subtended by each oriented boundary component can be summed to compute a generalized winding number of the entire shape (see Figure 3).
@@
For a watertight shape, the generalized winding number is a scalar field that partitions space into integer-valued regions which define containment in the shape. For non-watertight shapes, the values of this scalar field degrades away from integer values near geometric discontinuities and/or overlaps between boundary components. This degradation occurs smoothly, yet rapidly, and so the resulting real-valued field can be interpreted as defining containment in the same way, such as through rounding the scalar field to the nearest integer. Although other uses of the GWN field are possible, we are primarily concerned with its evaluation for the purposes of defining this robust containment query.
@@
2.2. Generalized Winding Numbers for Discretized Surfaces
The generalized winding number for open shapes was first conceptualized to efficiently evaluate containment queries on messy STL geometry, doing so via a hierarchical spatial index and divide-and-conquer algorithm (Jacobson et al., 2013). This work was later extended to entirely disconnected triangle soups by Barill et al. (2018), improving performance with an error-controlled approximation of the GWN with respect to groups of mesh elements that are far from the query point.
Evaluation methods for the GWN field defined by triangle meshes and soups are particularly computationally efficient, as the solid angle which defines the 3D GWN can be directly and exactly computed for polygonal shapes. However, if the input surface geometry is curved, a naïve application of these methods necessarily uses some geometric approximation of the surface to obtain the necessary input type, which in turn causes the GWN field of the approximation to diverge from the ground truth. At the same time, even generating a tessellation of a NURBS surface is a non-trivial task, and high-resolution meshes can be prohibitively expensive to generate and store for complex models (Xiong et al., 2023).
Another prominent GWN method comes from Barill et al. (2018), which defines the GWN of oriented point cloud data that could be sampled directly from a CAD-derived surface. Indeed, Balu et al. (2023) do so in the context of multiphysics simulation, describing a method which uses a point cloud randomly sampled from a CAD model to define an interior volume for the otherwise messy or complex shape. However, even if the weights for each point are accurately computed from the ground-truth surface (which may not be the case depending on implementation) discretizing the model into a finite collection of points results in significant errors in the evaluated GWN field, particularly for points near the surface.
@@ 2.2. Generalized Winding Numbers for Discretized Surfaces
The second is numerical imprecision in the computed fractional component of the GWN field. In 2D it is necessarily true that for a fixed set of input queries there exists some linear discretization of the input curve that produces the correct GWN field at each query, but this is not the case in 3D. Even if a query point is correctly placed on the proper side of the surface, the mismatch between the boundaries of the original surface and any triangulation introduces imprecision in the computed field (see Figure 4). While these errors decrease with increasing mesh resolution, they never vanish completely.
@@
2.3. Generalized Winding Numbers for General Surfaces
@@
The 2D GWN $w_{C}(q)$ at a query point $q$ with respect to a curve $C$ is defined as the total signed angle subtended by the curve, translated to place $q$ at the origin, normalized by a factor of $2\pi$. This results in the following integral of the (signed) differential angle $d\theta$,
@@
where CqC-q denotes translating the curve to make $q$ the origin. Written in Cartesian coordinates, this is equivalent to:
(2) $w_{C}(q)$ $=\frac{1}{2\pi}\int_{C-q}\frac{x\cdot\hat{n}}{\lVert x\rVert^{2}}\,dx.$
Evaluating this line integral directly via numerical quadrature is difficult, as the near-singular behavior of the integrand means that for points close to the curve, conventional quadrature methods exhibit severe instabilities. For linear elements with endpoints $a$ and $b$, however, this subtended angle has a simple and direct formula,
(3) $w_{L}:=\frac{1}{2\pi}\text{arctan}\left(\frac{\lVert a\times b\rVert}{a\cdot b}\right).$
@@
There exist a number of quadrature techniques in the literature that could be immediately applied to Equation 5, many of which can even be used in the presence of complicated trimming geometry (Gunderman et al., 2020; Chin and Sukumar, 2021; Gunderman et al., 2021). However, the integrand’s fundamental near-singularity makes direct numerical evaluation difficult for query points near the surface.
The hope, then, is that this too can be avoided through an exact formula for the analogous linear element. Indeed, the GWN for a 3D triangle defined by vertices $a$, $b$, and $c$ is given by
(6) $w_{\triangle}:=\frac{1}{2\pi}\text{arctan}\left(\frac{\scriptstyle\lVert a\cdot(b\times c)\rVert}{\scriptstyle\lVert a\rVert\lVert b\rVert\lVert c\rVert+(a\cdot b)\lVert c\rVert+(b\cdot c)\lVert a\rVert+(c\cdot a)\lVert b\rVert}\right).$
@@
To find a more direct and consistent solution method, we consider that in certain contexts, the 3D solid angle from which the GWN field is derived can be defined entirely from the boundary of the surface (Binysh and Alexander, 2018; Gillespie et al., 2024). In this spirit, one could imagine adapting the method of Jacobson et al. (2013) to a 3D curved surface $S$ by constructing a linear approximation of the surface boundary $\partial S^{\prime}$ (with $\partial S^{\prime}~\approx~\partial S$), along with a coarse triangulated closure $\overline{S^{\prime}}$ whose GWN field can be computed exactly. For query points far from the original surface, this approach is somewhat reasonable—although some precision is lost in the discretization of $\partial S^{\prime}$, the GWN field of the approximated closure is related to the GWN field of the original surface for far-away points through the simple formula $w_{S}~=~-w_{\overline{S}}~\approx~-w_{\overline{S^{\prime}}}$.
However, the simplicity and practicality of this boundary-focused approach necessarily breaks down for near-surface points. To reiterate, deriving $w_{S}$ from $w_{\overline{S}}$ in general requires calculating $w_{S\cup\overline{S}}$, i.e. the integer winding number of the query within the closed surface, which requires consideration of the original, undiscretized surface to preserve accuracy close to the surface. In this case, using $\overline{S^{\prime}}$ as a computational proxy for $\overline{S}$ complicates things further, as $w_{S\cup\overline{S^{\prime}}}$ is no longer integer valued, since $\overline{S}$ and $\overline{S^{\prime}}$ do not share the same boundary by construction.
@@
2.4. Generalized Winding Numbers for Trimmed NURBS Surfaces
@@
To explore the implications of this problem context, we consider the recent “One-Shot” approach of Martens and Bessmeltsev (2025) which, like the idealized adaptation of Jacobson et al. (2013) to curved surfaces in the previous section, evaluates the GWN of parametric surfaces with open edges by discretizing only the open edges within connected components of the shape. These polyline edges are projected onto a unit sphere around the query point, partitioning it into a graph of spherical polygons. After mapping this collection to integer crossing numbers determined through a single ray-surface intersection test, the total GWN is evaluated as the sum of each integer, weighted by the approximate surface area of the associated polygon. The quality of the boundary discretization influences both the precision of the area calculation and the accuracy of the crossing number assignment, even if the intersection test is performed with respect to the undiscretized surface.
Although the One-Shot approach is aimed at computing the GWN for parametric surfaces, we find its high-level approach unsuitable for our problem context. For example, a stated limitation of the One-Shot method is that it is most effective for surfaces with simple boundaries, as is typical for the preprocessed curve networks used as examples in that work. On the other hand, a trimmed NURBS surface can have an arbitrarily complicated boundary, as its projection onto the unit sphere can be combinatorially difficult to partition into distinct regions. Even so, as we demonstrate in Section 4.1, our method is more performant even when applied to simple shapes.
@@
3. Methods
In this work, we present a generalized winding number algorithm that can be accurately evaluated at arbitrary points in space relative to the CAD model, with precision up to an arbitrary user tolerance. We assume that each component surface is a trimmed tensor-product NURBS patch, with trimming curves that are explicitly defined in the parameter space of the knot spans of the patches by closed loops of 2D NURBS curves. This is the default representation for trimming curves in the STEP and IGES interchange formats and a common assumption in the literature (Piegl and Tiller, 1997; Sheng and Hirsch, 1992; Hamann and Tsai, 1996; Piegl and Richard, 1995; Sloup and Havran, 2021; Krishnamurthy et al., 2009). Since we use the 2D GWN method of Spainhour et al. (2024) to robustly determine containment in the 2D parameter space of each patch, we treat the collection of trimming curves as unstructured, which significantly simplifies their processing. For each query point, we evaluate our algorithm for each component surface in the model and sum the results, making the algorithm indifferent to watertightness by the linearity of the GWN.

### v3

@@ Abstract.
We propose a containment query that is robust to the watertightness of regions bound by trimmed NURBS surfaces, as this property is difficult to guarantee for in-the-wild CAD models. Containment is determined through the generalized winding number (GWN), a mathematical construction that is indifferent to the arrangement of surfaces in the shape. Applying contemporary techniques for the 3D GWN to trimmed NURBS surfaces requires some form of geometric discretization, introducing computational inefficiency to the algorithm and even risking containment misclassifications near the surface. In contrast, our proposed method leverages properties of the 3D solid angle to solve the relevant surface integral using a boundary formulation with rapidly converging adaptive quadrature. Batches of queries are further accelerated by memoizing (i.e. caching and reusing) quadrature node positions and tangents as they are evaluated. We demonstrate that our GWN method is robust to complex trimming geometry in a CAD model, and is accurate up to arbitrary precision at arbitrary distances from the surface. The derived containment query is therefore robust to model non-watertightness while respecting all curved features of the input shape.
@@
Graphical Abstract In the top half of the figure a machine part CAD shape with open edges is shown alongside a 2D slice through the middle showing the GWN field. The scalar field is rounded to the nearest integer to produce a plot of the containment decision. Two inset images are shown, one around a small open edge and one against a large open edge. In both cases, the containment decision shows a reasonable partition between inside and outside. In the bottom half of the figure the dimension-reduction properties of Stokes’ theorem are shown as divided into the three cases: “Far-Field”, “Near-Field”, and “Edge-Cases”.
@@ 1. Introduction
In computer graphics and adjacent fields, 3D objects are often defined in terms of a boundary representation (B-Rep) of the surface. While B-reps are very useful for flexibly designing 3D objects within a CAD system, they often fail to be watertight, in the sense that they do not define a legitimate interior volume (see Figure 2). For example, disconnected edges between component surfaces may be an intentional design decision, but are just as likely to result from human error during modeling, or differing tolerances between applications (Kasik et al., 2005; McAdams et al., 2011) (see Figure 2). The use of Boolean operations during model construction can also introduce problematic non-manifold edges. Issues of watertightness are greatly exacerbated in the context of trimmed NURBS (non-uniform rational B-spline) surfaces, where such gaps are nearly mathematically inevitable due to the complexities of projecting 3D surface-surface intersection curves into 2D parametric trimming curves (Sederberg et al., 2008; Zou et al., 2024). Nevertheless, we are often interested in determining whether a given point is contained within the shape in a way that conforms to the provided boundaries, but is robust enough to provide reasonable decisions in the vicinity of such geometric errors.
@@
Similarly, methods for immersogeometric analysis model fluid-structure interactions by placing a complicated CAD model on a simpler computational grid which records the numerical solution (Kamensky et al., 2015; Fromm et al., 2023). In such methods, it is necessary to know whether points on this grid are inside or outside the CAD model, which can be particularly challenging for points located arbitrarily close to the B-Rep.
@@
  • We extend the theory of generalized winding numbers to the context of trimmed NURBS patches. With this, we define a containment query for B-Reps composed of unstructured collections of such shapes that is robust to (and indeed indifferent to) watertightness.
  • We describe a novel algorithmic framework for evaluating the GWN with respect to a trimmed NURBS surface, greatly improving efficiency and stability by reformulating the relevant surface integral as a line integral along the surface boundary. We evaluate these integrals with an adaptive quadrature scheme to achieve results up to arbitrary accuracy, and further improve performance with a memoization strategy that caches and reuses geometric calculations for each surface. Within our algorithm, we robustly handle trim tests (visibility queries over trimming curves in the 2D parameter space of each surface) with the 2D GWN.
@@ 2.1. Containment Queries in 3D CAD Applications
There are many strategies to evaluate containment queries for CAD applications. In the typical ray casting method, an arbitrary ray is extended out from the point of interest and the number of intersections with bounding geometry are recorded. Assuming a watertight B-Rep, each such intersection increments a crossing number whose parity determines containment.
@@
There is considerable overlap with the adjacent task of ray tracing for the purposes of rendering, which typically considers a specific, pre-selected ray direction, and only requires calculation of the first intersection of a ray with the B-Rep. Much of the research focus in ray tracing relates to triangulated surfaces, as many hardware platforms are capable of calculating ray-triangle intersections directly, exactly, and massively in parallel (Haines and Akenine-Möller, 2019, Chapter 3). However, there is also interest in applying such techniques directly to trimmed NURBS surfaces (Schollmeyer and Fröhlich, 2009, 2019; Xiong et al., 2023; Zhu et al., 2025). Of particular note are recent methods for ray tracing bilinear patches, as, similar to triangles, they admit direct geometric formulae for computing their intersections with rays (Reshetov, 2019).
On more general curved surfaces, surface-line intersection routines are more complex and iterative (Nishita et al., 1990; Martin et al., 2000; Efremov et al., 2005; Shen et al., 2016; Marschner et al., 2021). Although the direction of the cast ray is arbitrary in principle, containment strategies for watertight shapes which utilize ray casting are highly sensitive to the specific configuration of surface and cast ray. This issue is compounded in the non-watertight case, as a cast ray passing through a numerical gap between surfaces results will result in an unexpectedly incorrect containment result.
Alternatively, winding number methods count the number of times the shape encloses the point, thereby implicitly handling configurations which are difficult for ray casting, even for watertight shapes (Carvalho and Cavalcanti, 1995; Hormann and Agathos, 2001). These methods can then be generalized by summing the signed angle subtended by individual boundary components so that the resulting containment decision is indifferent to watertightness (see Figure 3).
@@
The generalized winding number (GWN) is a real-valued scalar field that degrades smoothly, yet rapidly away from integer values near gaps and/or overlaps between boundary components. This field can be interpreted as defining containment, for example, by rounding its values to the nearest integer and applying either a non-zero or even-odd rule. Although other uses of the GWN field are possible, we are primarily concerned with its evaluation for the purposes of defining this robust containment query.
@@
Simple comparison of 2D and 3D GWN The GWN field is shown for two representations of a mechanical nut shape. On the top is a 2D representation, the 2D GWN in a bounding box of the shape, and an image depicting the GWN as a subtended angle. On the bottom is a 3D representation, a 2D slice of the 3D GWN field, and an image depicting the GWN as a subtended solid angle.
2.2. GWN Evaluation for Discretized Surfaces
The generalized winding number was first conceptualized to provide containment queries on messy STL geometry, doing so via a hierarchical spatial index and divide-and-conquer algorithm (Jacobson et al., 2013). This work was later extended to entirely disconnected triangle soups by Barill et al. (2018), improving computational performance with an error-controlled approximation of the GWN with respect to groups of triangles that are far from the query point.
Evaluation methods for the GWN field defined by triangle meshes and soups are particularly computationally efficient, as the solid angle which defines the 3D GWN can be directly and exactly computed for polygonal shapes. However, if the input surfaces are curved, a naïve application of these methods must necessarily use some geometric approximation of the surface to obtain the necessary input type, which in turn causes the GWN field of the approximation to diverge from the ground truth. At the same time, even generating a tessellation of a trimmed NURBS surface is a non-trivial task, and high-resolution meshes can be prohibitively expensive to generate and store for complex models (Xiong et al., 2023)
The work of Barill et al. (2018) defined the GWN of oriented point cloud data, which could be sampled directly from a CAD-derived surface. Indeed, Balu et al. (2023) do so in the context of multiphysics simulation, describing a method which uses a point cloud randomly sampled from a CAD model to define an interior volume for the otherwise messy or complex shape. However, even if the weights for each point are accurately computed from the ground-truth surface (which may not be the case depending on implementation), such a discretization results in significant errors in the evaluated GWN field, particularly for points near the surface.
@@ 2.2. GWN Evaluation for Discretized Surfaces
The second is numerical imprecision in the computed fractional component of the GWN field. While it is necessarily true in 2D that, for a fixed set of input queries, there exists some linear discretization of the input curve that produces the correct GWN field at each query, this is no longer the case in 3D. Even if a query point is correctly placed on the proper side of the surface, the mismatch between the boundaries of the original surface and any triangulation introduces imprecision in the computed field that is difficult to predict from the triangulation alone (see Figure 4). While these errors decrease with increasing mesh resolution, they never vanish completely.
@@
Comparison between the GWN of a curved/discretized object in 2D/3D On the top is shown the GWN field for a 2D curve, a polyline which approximates the curve, and a plot of the absolute error between them. It is shown that the error is either 0 or 1 everywhere. On the bottom is the GWN for a 3D surface, a triangulation which approximates the surface, and a plot of the absolute error between them. It is shown that the error is nonzero everywhere.
@@
2.3. GWN Evaluation for General Surfaces
@@
The 2D GWN $w_{C}(q)$ at a query point $q$ with respect to a curve $C$ is defined as the total signed angle subtended by the curve, translated to place $q$ at the origin, and normalized by a factor of $2\pi$. This results in the following integral of the (signed) differential angle $d\theta$,
@@
where CqC-q denotes translating the curve to make $q$ the origin.
Evaluating this line integral directly via numerical quadrature is difficult, as the near-singular behavior of the integrand means that for points close to the curve, conventional quadrature methods are very unstable. For linear elements with (translated) endpoints $a$ and $b$, however, this subtended angle has a simple and direct formula,
(2) $w_{L-q}(0):=\frac{1}{2\pi}\text{arctan}\left(\frac{\lVert a\times b\rVert}{a\cdot b}\right).$
@@
There exist a number of quadrature techniques in the literature that could be immediately applied to Equation 4, many of which can in principle be used even in the presence of complicated trimming geometry (Gunderman et al., 2020; Chin and Sukumar, 2021; Gunderman et al., 2021). However, the integrand’s fundamental near-singularity makes direct numerical evaluation difficult for query points near the surface.
The hope, then, is that this too can be avoided through an exact formula for the analogous linear element. Indeed, the GWN for a 3D triangle defined by (translated) vertices $a$, $b$, and $c$ is given by
(5) $w_{\triangle-q}(0):=\frac{1}{2\pi}\text{arctan}\left(\frac{\scriptstyle\lVert a\cdot(b\times c)\rVert}{\scriptstyle\lVert a\rVert\lVert b\rVert\lVert c\rVert+(a\cdot b)\lVert c\rVert+(b\cdot c)\lVert a\rVert+(c\cdot a)\lVert b\rVert}\right).$
@@
To find a more direct and consistent solution method, we utilize a known property of the 3D solid angle which states that in certain contexts, it and the derived GWN field can be defined entirely from the boundary of the surface (Jacobson et al., 2013; Binysh and Alexander, 2018; Gillespie et al., 2024). In this spirit, one could imagine adapting the method of Jacobson et al. (2013) to a 3D curved surface $S$ by constructing a linear approximation of the surface boundary $\partial S^{\prime}$ (with $\partial S^{\prime}~\approx~\partial S$), along with a coarse triangulated closure $\overline{S^{\prime}}$ whose GWN field can be computed exactly. For query points far from the original surface, this approach is somewhat reasonable—although some precision is lost in the discretization of $\partial S^{\prime}$, the GWN field of the approximated closure is related to the GWN field of the original surface for far-away points through the simple formula $w_{S}~=~-w_{\overline{S}}~\approx~-w_{\overline{S^{\prime}}}$.
However, the simplicity and practicality of this boundary-focused approach necessarily breaks down for near-surface points. To reiterate, deriving $w_{S}$ from $w_{\overline{S}}$ in general requires calculating $w_{S\cup\overline{S}}$, i.e. the integer winding number of the query within the closed surface, which requires consideration of the original, undiscretized surface to preserve accuracy close to the surface. In this case, using $\overline{S^{\prime}}$ as a computational proxy for $\overline{S}$ complicates things further, as $w_{S\cup\overline{S^{\prime}}}$ is no longer integer valued, since, by construction, $\overline{S}$ and $\overline{S^{\prime}}$ do not share the same boundary.
@@
2.4. GWN Evaluation for Trimmed NURBS Surfaces
@@
To explore the implications of this problem context, we consider the recent “One-Shot” approach of Martens and Bessmeltsev (2025) which, like the idealized adaptation of Jacobson et al. (2013) to curved surfaces in the previous section, evaluates the GWN of parametric surfaces with open edges by discretizing only the open edges within connected components of the shape. These polyline edges are projected onto a unit sphere around the query point, partitioning it into a graph of spherical polygons. After mapping this collection to integer crossing numbers determined through a single ray-surface intersection test, the total GWN is evaluated as the sum of each integer, weighted by the approximate surface area of the associated polygon. The quality of the boundary discretization naturally influences the precision of the area calculation, but also the accuracy of the subsequent containment decision.
While the suggested ray-surface intersection routine does take into consideration the original (undiscretized) surface, the mapping of crossing numbers to spherical polygons may still introduce off-by-one errors depending on the quality of the boundary discretization. Although careful selection of the ray direction may reduce the likelihood of misclassification, this risk represents a fundamental limitation of using unconsidered discretizations of the input shape.
Although the One-Shot approach is aimed at computing the GWN for parametric surfaces, we find its high-level approach unsuitable for our problem context. For example, a stated limitation of the One-Shot method is that it is most effective for surfaces with simple boundaries, as is typical for the preprocessed curve networks used as examples in that work. On the other hand, a trimmed NURBS surface can have an arbitrarily complicated boundary, and so partitioning the unit sphere projection of these boundaries becomes combinatorially intractable in practical contexts. Importantly, while the cost of the line-surface intersection procedure is reduced for query points which are far from the surface (and could theoretically be eliminated with minor algorithmic adjustments), the cost of this partitioning procedure remains relatively constant across space, and even demands a higher-resolution boundary discretization as query points become further away and partitioned regions become smaller. Even so, as we demonstrate in Section 4.1, our method is more performant even when applied to simple shapes.
@@
3. Method
In this work, we present a generalized winding number algorithm that can be accurately evaluated at arbitrary points in space relative to the CAD model, with precision up to an arbitrary user tolerance. We assume that each component surface is an oriented trimmed tensor-product NURBS patch, with trimming curves that are explicitly defined in the parameter space of the knot spans of the patches by closed loops of 2D NURBS curves. This is the default representation for trimming curves in the STEP and IGES interchange formats and a common assumption in the literature (Piegl and Tiller, 1997; Sheng and Hirsch, 1992; Hamann and Tsai, 1996; Piegl and Richard, 1995; Sloup and Havran, 2021; Krishnamurthy et al., 2009). The orientation of the patch is determined by the direction of each parameter-space knot span. Since we use the 2D GWN method of Spainhour et al. (2024) to robustly determine containment in the parameter space of each patch, we treat the collection of trimming curves as unstructured, significantly simplifying their processing. For each query point, we evaluate our algorithm for each component surface in the model and sum the results, making the algorithm indifferent to watertightness by the linearity of the GWN.